OpenStudy (anonymous):

The harvest increased by 120%. If the harvest is 125 bushels, what was the harvest last year? a. 150 b. 96 c. 56.8 d. 100 e. none of the above

jimthompson5910 (jim_thompson5910):

x = harvest last year

jimthompson5910 (jim_thompson5910):

x + (120% of x) = 125 x + 1.2x = 125 2.2x = 125 x = 125/2.2 x = 56.8181818181819 x = 56.8 So the harvest last year was 56.8 bushels